OpenStudy (anonymous):

geometry/trigonometry question medal, fan, and testimonial *i need someone to teach me how to do this. I'm pretty sure if this had numbers I'd be able to solve it. btw if you plan to teach me pleaseee be patient and try to teach me as if you're teaching a 3rd grader lol :] thanks*

OpenStudy (anonymous):

i can help you with this

OpenStudy (anonymous):

Okay thank you :]

OpenStudy (anonymous):

the trig identities can easily be remembered using the mnemonic soh-cah-toa. sine: opposite/hypotenuse cosine: adjacent/hypotenuse tangent: opposite/adjacent

OpenStudy (anonymous):

so when you have sin(a)= (side opposite of angle a)/(hypotenuse) or sin(a)=x/z

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

Yeah I know what they are and yes that makes sense.

OpenStudy (anonymous):

ok so then try b and let me know what you come up with

OpenStudy (anonymous):

Wait hold on so.. the answer is just the "equation" as I like to call it. sin(a)=x/z? though I thought you put it like sin(a)x/z I didn't know you had to put an equal sign.

OpenStudy (anonymous):

yes the answer will be an equation since there are no numbers given. sin(a)=x/z. no the equals sign must be there :)

OpenStudy (anonymous):

Okay cool I thought I had to solve something xD I can't stop laughing now but anyway I'll do the second one and tell me if I did right okay? :]

OpenStudy (anonymous):

nope, just the equation :) its alright it happens to all of us! yes go ahead and do that and ill let you know if its right :)

OpenStudy (anonymous):

cos(a)=y/x ?

OpenStudy (anonymous):

close, but be careful with the hypotenuse, z is the hypotenuse not x. it would be cos(a)=y/z

OpenStudy (anonymous):

Ohhh oops yeah I see now lol. Okay well thank you so much!! I think I got it from here :] thank youuuuu #lifesaver

OpenStudy (anonymous):

yeah easy mistake to make its ok :) alright good luck! if you want to type your answers here when youre done i could check them for you! my pleasure to help! :)

OpenStudy (anonymous):

Thank you sure :] also thanks for being nice. Sometimes people loose their patience with me lol.

OpenStudy (anonymous):

Yeah you bet, you were easy to teach :)

OpenStudy (anonymous):

Hehe thanks. I honestly feel like people here on openstudy have better luck at teaching me than my teachers do xD honestly that says a lot about the school system.